# A classification of Fourier summation formulas and crystalline measures

### v1

@@ 7. Proof of Theorem 1
is the Dedekind eta-function, $q=e^{2\pi iz}$ and $\chi_{12}$ is the Dirichlet character of modulus $12$ with $\chi(1)=\chi(11)=-\chi(5)=-\chi(7)=1$ and zero otherwise. It is clear that $\operatorname{\mathrm{supp}}(\mu)$ intersects any infinite arithmetic progression at most finitely.
We now generalize Guinand’s example broadly. For that we need the following lemma.
@@
Let $F\in\mathrm{AP}(\mathbb{C}^{+})$, with locally finite spectrum contained in $[0,\infty)$ and assume that
@@
for some $k>0$, where ${\rm spec}(F)=\{\gamma_{n}\}_{n\geq 0}$, Assume that $G(z)=F(-1/z)\sqrt{i/z}$ also belongs to AP(+)normal-$\mathrm{AP}(\mathbb{C}^{+})$, has locally finite spectrum contained in $[0,\infty)$ and $\mathbb{E}G$ decays similarly as $\mathbb{E}F$. Let
$\mu=\sum_{n\geq 0}\mathbb{E}F(\gamma_{n})({\boldsymbol{\delta}}_{-\sqrt{2\gamma_{n}}}+{\boldsymbol{\delta}}_{-\sqrt{2\gamma_{n}}}).$
@@
The proof is inspired by [13, Lemma 2.1]. Consider the gaussian $g_{z}(t)=e^{\pi izt^{2}}$ for $z\in\mathbb{C}^{+}$. Observe that $\widehat{g}_{z}(t)=\sqrt{i/z}g_{-1/z}(t)$ and that $g_{z}(\sqrt{2s})=e^{2\pi izs}$ for $s\geq 0$. Assume first that $\mathbb{E}G$ also has polynomial growth. Then the identity
$G(z)=\sum_{n\geq 0}\mathbb{E}G(\lambda_{n})g_{z}(\sqrt{2\lambda_{n}})=F(-1/z)\sqrt{i/z}=\sum_{n\geq 0}\mathbb{E}F(\gamma_{n})\widehat{g}_{z}(\sqrt{2\lambda_{n}})$
@@
$\sum_{n\geq 0}\mathbb{E}F(\gamma_{n})(\widehat{\varphi}(\sqrt{2\gamma_{n}})+\widehat{\varphi}(\sqrt{-2\gamma_{n}}))=\sum_{n\geq 0}\mathbb{E}G(\lambda_{n})(\varphi(\sqrt{2\lambda_{n}})+\varphi(-\sqrt{2\lambda_{n}})),$
for any $\varphi\in{\rm span}_{\mathbb{C}}\{g_{z}:z\in\mathbb{C}^{+}\}$. An approximation argument similar to the one employed in [13, Lemma 2.1] shows that the above identity actually holds for any $\varphi\in\mathcal{S}(\mathbb{R})$. ∎
The following construction was inspired by a similar one partially communicated in 2018 by Danylo Radchenko in a seminar at the University of Bonn. Let $N\geq 1$ be an integer and ${\boldsymbol{r}}=\{r_{d}\}_{d|N}\subset\mathbb{Q}$ be a sequence of rationals indexed by the divisors of $N$ such that
$r_{d}=r_{N/d},\quad\sum_{d|N}r_{d}=1\ \text{ and }\ \sum_{d|N}dr_{d}=\frac{24k}{b}$ (19)
where $b\geq 1$ and $k\geq 0$ are coprime integers with $b=1$ if $k=0$. Consider the eta-product
$\eta({\boldsymbol{r}},z)=\prod_{d|N}\eta(dz)^{r_{d}}=\sum_{n\geq k}\alpha_{n}q^{n/b}.$
The eta-function $\eta:\mathbb{C}^{+}\to\mathbb{C}$ is an holomorphic function that satisfies the functional identities $\eta(z+\ell)=e^{\pi i\ell/12}\eta(z)$, for $\ell\in\mathbb{Z}$, and $\eta(-1/z)=\sqrt{z/i}\,\eta(z)$. We obtain that
$\eta({\boldsymbol{r}},z+b)=e^{\pi ib(\sum_{d|N}dr_{d})/12}\eta({\boldsymbol{r}},z)=\eta({\boldsymbol{r}},z)$
@@
Above we use that $\prod_{d|N}d^{-r_{d}/2}=N^{-1/4}$, which follows from the conditions on $r_{d}$. In particular we conclude that the function
$F_{+}({\boldsymbol{r}},z)=\eta({\boldsymbol{r}},z/\sqrt{N})=\sum_{n\geq k}\alpha_{n}q^{n/(b\sqrt{N})}.$
@@
$F_{+}({\boldsymbol{r}},z+b\sqrt{N})=F_{+}({\boldsymbol{r}},z)\quad\text{and}\quad\sqrt{i/z}\,F_{+}({\boldsymbol{r}},-1/z)=F_{+}({\boldsymbol{r}},z)$
for all $z\in\mathbb{C}^{+}$. It can be shown that the function $\eta({\boldsymbol{r}},z)$ is a holomorphic modular form of level $N$, and so the Hecke bound produces $|\alpha_{n}|\ll n^{1/4}$ (see also [8]). Hence $F_{+}$ satisfies the hypotheses of Lemma 19. Thus, the measure
$\mu_{+}=\sum_{n\geq k}\alpha_{n}({\boldsymbol{\delta}}_{\sqrt{2n/(b\sqrt{N})}}+{\boldsymbol{\delta}}_{-\sqrt{2n/(b\sqrt{N})}}),$
satisfies
$\widehat{\mu}_{+}=\mu_{+}.$
In particular, $\mu_{+}$ is strongly tempered. Poisson Summation can be recovered this way by
$\frac{\eta(2z)^{5}}{\eta(z)^{2}\eta(4z)^{2}}=1+2q+2q^{4}+2q^{9}+2q^{25}+\cdots,$
with parameters $N=4$, $(r_{1},r_{2},r_{4})=(-2,5,-2)$, $b=1$ and $k=0$. Guinand’s example (18) has parameters $N=4$, $(r_{1},r_{2},r_{4})=(2/3,-1/3,2/3)$, with the others $r_{d}=0$ for $d|36$, $b=3$ and $k=1$.
Assume now that $N$ is a square. Using the modular $\lambda$-invariant (Hauptmodul for $\Gamma(2)$)
$\lambda(z)=\frac{16\eta(2z)^{16}\eta(z/2)^{8}}{\eta(z)^{24}}=16q^{1/2}-128q+704q^{3/2}+\cdots,$
which can be shown to satisfy the transformation laws
$\lambda(-1/z)=1-\lambda(z)\quad,\text{and}\quad\lambda(z+1)=\lambda(z)/(\lambda(z)-1),$
we can define instead
$F_{-}({\boldsymbol{r}},z)=(1-2\lambda(z))\eta({\boldsymbol{r}},z/\sqrt{N})=\sum_{n\geq 2k}\beta_{n}q^{n/(2b\sqrt{N})}$
and deduce that
$F_{-}({\boldsymbol{r}},z+2b\sqrt{N})=F_{-}({\boldsymbol{r}},z)\quad\text{ and }\quad\sqrt{i/z}\,F_{-}({\boldsymbol{r}},-1/z)=-F_{-}({\boldsymbol{r}},z).$
Again we conclude that the measure
$\mu_{-}=\sum_{n\geq 2k}\beta_{n}({\boldsymbol{\delta}}_{\sqrt{n/(b\sqrt{N})}}+{\boldsymbol{\delta}}_{-\sqrt{n/(b\sqrt{N})}}),$
satisfies $\widehat{\mu}_{-}=-\mu_{-}.$
It is now obvious that one can build infinitely many examples of such measures $\mu_{\pm}=\pm\widehat{\mu}_{\pm}$ with locally finite support. An interesting family of examples is
$\frac{\eta(z)^{l}\eta(4z)^{l}}{\eta(2z)^{2l-1}}$ $=q^{(l+2)/24}\left(1-lq+\frac{(l-1)(2+l)}{2}q^{2}+\sum_{n\geq 3}\alpha_{n,l}q^{n}\right)$
$(1-2\lambda(2z))\frac{\eta(z)^{l}\eta(4z)^{l}}{\eta(2z)^{2l-1}}$ $=q^{(l+2)/24}\left(1-(32+l)q+\tfrac{1}{2}(l^{2}+65l+510)q^{2}+\sum_{n\geq 3}\beta_{n,l}q^{n}\right)$
for $N=4$ and rational $l\geq-2$. It is not hard to show that $n!\alpha_{n,l}$ and $n!\beta_{n,l}$ are polynomials in $l$ with integer coefficients and degree at most $n$. The examples above produce measures $\mu_{\pm}=\pm\widehat{\mu}_{\pm}$ with support contained in
$\{\sqrt{c+n}\}_{n\geq 0}\cup\{-\sqrt{c+n}\}_{n\geq 0},$
for any rational $c=(l+2)/24\geq 0$. Thus, for non integers $c$, the supports of $\mu_{\pm}$ intersect any infinite arithmetic progression at most finitely. Note that for $l=-2$ we recover classical Poisson summation and for $l=2/3$ we recover Guinand’s example. Other interesting constructions may be derived from [24] and the references therein.
@@
Such measures are also connected with new Fourier interpolation formulae produced in [33], which as a byproduct produces crystalline pairs with space and spectral supports contained in {±$\{\pm\sqrt{n}\}_{n\geq 0}\cup\{x_{0}\}$, for any $x_{0}$ not an integer square root. The recent papers [34, 22] also suggest that there should exist a crystalline pair with space and spectral supports contained in {±$\{\pm n^{\alpha}\}_{n\geq 1}$ and {±$\{\pm n^{(1-\alpha)}\}_{n\geq 1}$ respectively, whenever $10<\alpha<$ 1. Moreover, [22, Cor. 3] proves the striking result that for every (supercritical) pair of increasing sequences $\mathcal{T}=\{t_{j}\}_{j\in\mathbb{Z}}$ and $\mathcal{S}=\{s_{j}\}_{j\in\mathbb{Z}}$ such that
@@
for some $1<p,q<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, there is a $\mathrm{FS}$-pair $(\mu,a)$ with $\mu$ supported in $\mathcal{T}$ and $a(\cdot)$ supported in $\mathcal{S}$. An important simple example is $\mathcal{T}=\{\pm n^{\alpha}\}_{n\geq 0}$ and $\mathcal{S}=\{\pm n^{\beta}\}_{n\geq 0}$ for positive $\alpha,\beta$ with $\alpha+\beta<1$. In fact, $(\mathcal{T},\mathcal{S})$ is both a Fourier uniqueness set and an interpolating set).
Almost Crystalline Measures
We say $(\mu,a)$ is an almost Fourier summation pair ($\mathrm{AFS}$-pair) if $\mu$ is strongly tempered, $a(\cdot)$ is locally summable function and $(a-\widehat{\mu})|_{C_{c}^{\infty}(\mathbb{R})}=(\nu_{1}-\widehat{\nu}_{2})|_{C_{c}^{\infty}(\mathbb{R})}$, where $\nu_{1}$ and $\nu_{2}$ are absolutely continuous, $\nu_{1}\neq 0$ is tempered and $\nu_{2}$ is strongly tempered. In particular, if $\varphi\in C_{c}^{\infty}(\mathbb{R})$ we have
$\sum_{\lambda\in\mathbb{R}}a(\lambda)\varphi(\lambda)-\int_{\mathbb{R}}\widehat{\varphi}(t)\mathrm{d}\mu(t)=\int_{\mathbb{R}}\varphi(t)\mathrm{d}\nu_{1}(t)-\int_{\mathbb{R}}\widehat{\varphi}(t)\mathrm{d}\nu_{2}(t).$
We then let the class $\rm ACM$ (Almost Crystalline Measures) denote the $\mathrm{AFS}$-pairs $(\mu,a)$ such that both $\mu$ and $a(\cdot)$ have locally finite support. A typical example is Guinand’s prime summation formula, which states that for every $\varphi\in C_{c}^{\infty}(\mathbb{R})$ we have
$\frac{\log(\pi/N)}{2\pi}\varphi(0)+\frac{1}{2\pi}\sum_{n\geq 1}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\bigg{(}\varphi\left(\frac{\log n}{2\pi}\right)+\varphi\left(\frac{-\log n}{2\pi}\right)\bigg{)}$
$=-\sum_{\gamma}\widehat{\varphi}(\gamma)+2\int_{\mathbb{R}}\varphi(t)\cosh(\pi t)\mathrm{d}t+\frac{1}{2\pi}\int_{\mathbb{R}}\widehat{\varphi}(t){\rm Re}\,\psi(\tfrac{1}{4}+i\tfrac{t}{2})\mathrm{d}t$
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is the digamma function, $\rho=1/2+i\gamma$ are the zeros of the Dirichlet $L$-function $L(\chi,s)=\sum_{n\geq 1}\chi(n)n^{-s}$ on the critical strip $10<{\rm Re}\,s<$ 1 and $\chi$ is a primitive, real and even Dirichlet character of modulus $N$. Since $|\psi(1/4+it/2)|\ll\log(2+|t|)$ and $\sum_{\gamma}\frac{1}{1+\gamma^{2}}<\infty$ we conclude, if we let assuming the Riemann Hypothesis for $L(\chi,s)$, that

### v3

@@ 8.4. Crystalline Measures ($\rm CM$)
is the Dedekind eta-function, $q=e^{2\pi iz}$ and $\chi_{12}$ is the Dirichlet character of modulus $12$ with $\chi_{12}(1)=\chi_{12}(11)=-\chi_{12}(5)=-\chi_{12}(7)=1$ and zero otherwise. Note that if $n=9m^{2}+2m$ then $\sqrt{n+1/9}=3m+1/3$, and so the support of $\mu$ contains an infinite arithmetic progression $\{3m+1/3\}_{m\geq 1}$. We will show below that, with the right point of view, one can embed Guinand’s example and also classical Poisson summation in a real one-parameter family, for which almost all members have the property that their support intersect any infinite arithmetic progression at most four times. For that we need first the following lemma.
@@
Let $F\in\mathrm{AP}(\mathbb{C}^{+})$, with locally finite spectrum contained in $0,\infty)$ and assume that
@@
for some $k>0$, where ${\rm spec}(F)=\{\gamma_{n}\}_{n\geq 0}$. Assume that $G(z)=F(-1/z)\sqrt{i/z}$ also belongs to $\mathrm{AP}(\mathbb{C}^{+})$, has locally finite spectrum contained in $0,\infty)$ and $\mathbb{E}G$ satisfies a similar estimate the one above for $\mathbb{E}F$. Let
$\mu=\sum_{n\geq 0}\mathbb{E}F(\gamma_{n})({\boldsymbol{\delta}}_{\sqrt{2\gamma_{n}}}+{\boldsymbol{\delta}}_{-\sqrt{2\gamma_{n}}}).$
@@
The proof is inspired by [17, Lemma 2.1]. Consider the gaussian $g_{z}(t)=e^{\pi izt^{2}}$ for $z\in\mathbb{C}^{+}$. Observe that $\widehat{g}_{z}(t)=\sqrt{i/z}g_{-1/z}(t)$ and that $g_{z}(\sqrt{2s})=e^{2\pi izs}$ for $s\geq 0$. Then the identity
$G(z)=\sum_{n\geq 0}\mathbb{E}G(\lambda_{n})g_{z}(\sqrt{2\lambda_{n}})=F(-1/z)\sqrt{i/z}=\sum_{n\geq 0}\mathbb{E}F(\gamma_{n})\widehat{g}_{z}(\sqrt{2\gamma_{n}})$
@@
$\sum_{n\geq 0}\mathbb{E}F(\gamma_{n})(\widehat{\varphi}(\sqrt{2\gamma_{n}})+\widehat{\varphi}(-\sqrt{2\gamma_{n}}))=\sum_{n\geq 0}\mathbb{E}G(\lambda_{n})(\varphi(\sqrt{2\lambda_{n}})+\varphi(-\sqrt{2\lambda_{n}})),$
for any $\varphi\in{\rm span}_{\mathbb{C}}\{g_{z}:z\in\mathbb{C}^{+}\}$. An approximation argument similar to the one employed in [17, Lemma 2.1] shows that the above identity actually holds for any $\varphi\in\mathcal{S}(\mathbb{R})$. ∎
The following constructions were inspired by a similar one partially communicated in 2018 by Danylo Radchenko in a seminar at the University of Bonn.
8.4.1. $N$-Level eta-quotients and self-dual crystalline measures
Let $N\geq 1$ be an integer and ${\boldsymbol{r}}=\{r_{d}\}_{d|N}$ be a sequence of reals indexed by the divisors of $N$ such that
$r_{d}=r_{N/d},\quad\sum_{d|N}r_{d}=1\ \text{ and }\ \frac{1}{24}\sum_{d|N}dr_{d}=b,$ (19)
where $c\geq 0$. Consider the eta-quotient
$\eta({\boldsymbol{r}},z)=\prod_{d|N}\eta(dz)^{r_{d}}=q^{b}\sum_{n\geq 0}\alpha_{n}q^{n}.$
The eta-function $\eta:\mathbb{C}^{+}\to\mathbb{C}$ is an holomorphic function that satisfies the functional identities $\eta(z+\ell)=e^{\pi i\ell/12}\eta(z)$, for $\ell\in\mathbb{Z}$, and $\eta(-1/z)=\sqrt{z/i}\,\eta(z)$. In particular we obtain that
$\eta({\boldsymbol{r}},z+1)=e^{2\pi ib}\eta({\boldsymbol{r}},z)$
@@ 8.4.1. $N$-Level eta-quotients and self-dual crystalline measures
Above we used that $\prod_{d|N}d^{-r_{d}/2}=N^{-1/4}$, which follows from the conditions on $r_{d}$. Eta-quotients of this type are known to be weakly holomorphic functions of weight $1/2$ on $\Gamma_{0}(N)$ (see [11, Prop. 1.39]). In particular, we conclude that the function
$F({\boldsymbol{r}},z)=\eta({\boldsymbol{r}},z/\sqrt{N})=q^{b/\sqrt{N}}\sum_{n\geq k}\alpha_{n}q^{n/\sqrt{N}}.$
@@
$F({\boldsymbol{r}},z+\sqrt{N})=e^{2\pi ib}F({\boldsymbol{r}},z)\quad\text{and}\quad\sqrt{i/z}\,F({\boldsymbol{r}},-1/z)=F({\boldsymbol{r}},z)$
for all $z\in\mathbb{C}^{+}$. Since the function $({\rm Im}\,z)^{1/4}|\eta({\boldsymbol{r}},z)|$ is $\Gamma_{0}(N)$-invariant, one can straightforwardly deduce that $|\alpha_{n}|\ll n^{k}$, for some $k>0$, whenever $\eta({\boldsymbol{r}},z)$ grows at most polynomially at all cusps of $\Gamma_{0}(N)$ (one can even obtain the Hecke-bound $|\alpha_{n}|\ll n^{1/4}$ if $\eta({\boldsymbol{r}},z)$ vanishes exponentially at all cusps of $\Gamma_{0}(N)$). This condition turns out to be equivalent to the verification of at most $d(N)$-linear inequalities given by [^20] We are thankful to D. Radchenko for pointing this out. [11, Lemma 1.40]. Assuming that such inequalities are satisfied, $F$ satisfies the hypotheses of Lemma 20, and the measure
@@
$\mu=\sum_{n\geq 0}\alpha_{n}({\boldsymbol{\delta}}_{\sqrt{2(n+b)/\sqrt{N}}}+{\boldsymbol{\delta}}_{-\sqrt{2(n+b)/\sqrt{N}}})\quad\text{satisfies }\quad\widehat{\mu}=\mu.$
It is now obvious that one can build infinitely many examples of self-dual crystalline measures $\mu$ by making linear combination of these types of eta-quotients.
For instance, consider now the following family
$\frac{\eta(z)^{24c-2}\eta(4z)^{24c-2}}{\eta(2z)^{48c-5}}$ $=q^{c}\left(1-(24c-2)q+(288c^{2}-36c)q^{2}+\sum_{n\geq 3}\alpha_{n,c}q^{n}\right),$
which fits the previous discussion for $N=4$, $(r_{1},r_{2},r_{4})=(24c-2,5-48c,24c-2)$ and $b=c$. We will have to impose $c\in[0,1/8]$, as this range guarantees that the coefficients satisfy that Hecke bound $|\alpha_{n,c}|\ll n^{1/4}$, since it will be holomorphic at the cusps 0, $i\infty$ and $1/2$ of $\Gamma_{0}(4)$ (actually, numerically, the coefficients $\alpha_{n,c}$ seem to be bounded). The example above produces a measure $\mu_{c}$ such that
$\mu_{c}=\widehat{\mu}_{c}\quad\text{and}\quad\operatorname{\mathrm{supp}}(\mu_{c})\subset\{\sqrt{c+n}\}_{n\geq 0}\cup\{-\sqrt{c+n}\}_{n\geq 0}.$
Note that $\mu_{0}$ is classical Poisson summation and $\mu_{\tfrac{1}{9}}$ is Guinand’s example. Perhaps other interesting constructions may be derived from [29] and the references therein. It is not hard to show that $\{\alpha_{n,c}\}_{n\geq 0}$ are polynomials in $c$ with integer coefficients and degree $n$. To finish the discussion, note that if $c$ is irrational then $\{\sqrt{c+n}\}_{n\geq 0}$ intersects any infinite arithmetic progression at most $2$ times. Indeed, having three hits implies that
$\frac{\sqrt{n_{2}+c}-\sqrt{n_{1}+c}}{k_{1}}=\frac{\sqrt{n_{3}+c}-\sqrt{n_{2}+c}}{k_{2}}$
is satisfied for some integers $n_{3}>n_{2}>n_{1}\geq 0$ and $k_{1},k_{2}>0$. Writing $A=n_{2}+c$, $n_{3}+c=A+k_{3}$, $n_{1}+c=A-k_{4}$ and solving for $A$, shows that $A$ is rational, which is absurd.
@@
Such measures are also connected with new Fourier interpolation formulae produced in [42], which as a byproduct produces crystalline pairs with space and spectral supports contained in $\{\pm\sqrt{n}\}_{n\geq 0}\cup\{x_{0}\}$, for any $x_{0}$ not an integer square root. These nodes can also be perturbed to be $\{\pm\sqrt{n+\varepsilon_{n}}\}_{n\geq 0}$, as long as $\varepsilon_{0}=0$ and $|\varepsilon_{n}|\leq\delta n^{-5/4}$ for some small $\delta>0$, see [44]. The recent papers [43, 27] also suggest that there should exist a crystalline pair with space and spectral supports contained in $\{\pm(\tfrac{1}{2\alpha}n)^{\alpha}\}_{n\geq 0}$ and $\{\pm(\tfrac{1}{2(1-\alpha)}n)^{1-\alpha}\}_{n\geq 0}$ respectively, whenever $0<\alpha<1$. Moreover, [27, Cor. 3] proves the striking result that for every (supercritical) pair of unbounded increasing sequences $\mathcal{T}=\{t_{j}\}_{j\in\mathbb{Z}}$ and $\mathcal{S}=\{s_{j}\}_{j\in\mathbb{Z}}$ such that
@@
for some $1<p,q<\infty$ with $\frac{1}{p}+\frac{1}{q}=1$, there is a $\mathrm{FS}$-pair $(\mu,a)$ with $\mu$ supported in $\mathcal{T}$ and $a(\cdot)$ supported in $\mathcal{S}$. An important simple example is again $\mathcal{T}=\{\pm(\tfrac{a}{2\alpha}n)^{\alpha}\}_{n\geq 0}$ and $\mathcal{S}=\{\pm(\tfrac{b}{2\beta}n)^{\beta}\}_{n\geq 0}$ for $0<\alpha,\beta,a,b<1$ with $\alpha+\beta\leq 1$ (the multiplying constants $a$ and $b$ are irrelevant unless $\alpha+\beta=1$, in which case we must impose $ab<1$, but conjecturally $ab=1$ is possible). In fact, $(\mathcal{T},\mathcal{S})$ is both a Fourier uniqueness pair and has an associated reconstruction/interpolation formula (for more info see [27]).
8.5. Almost Crystalline Measures
We say $(\mu,a)$ is an almost Fourier summation pair ($\mathrm{AFS}$-pair) if $\mu$ is strongly tempered, $a(\cdot)$ is locally summable function and
$(\widehat{\mu}-a)|_{C_{c}^{\infty}(\mathbb{R})}=\nu|_{C_{c}^{\infty}(\mathbb{R})},$
for some absolutely continuous complex-measure $\nu$. In particular, if $\varphi\in C_{c}^{\infty}(\mathbb{R})$ we have
$\sum_{\lambda\in\mathbb{R}}a(\lambda)\varphi(\lambda)+\int_{\mathbb{R}}\varphi(t)\mathrm{d}\nu(t)=\int_{\mathbb{R}}\widehat{\varphi}(t)\mathrm{d}\mu(t).$
One way to think about this that $\nu$ is the absolutely continuous part of $\widehat{\mu}$, while $a(\cdot)$ is the (atomic) singular part. We then let the class $\rm ACM$ (Almost Crystalline Measures) denote the $\mathrm{AFS}$-pairs $(\mu,a)$ such that both $\mu$ and $a(\cdot)$ have locally finite support. The typical example is Guinand’s prime summation formula (also known as Riemann–Weil explicit formula), which states that for every $\varphi\in C_{c}^{\infty}(\mathbb{R})$ we have
$\frac{\log(\pi/N)}{2\pi}\varphi(0)+\frac{1}{2\pi}\sum_{n\geq 2}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\bigg(\varphi\left(\frac{\log n}{2\pi}\right)+\varphi\left(\frac{-\log n}{2\pi}\right)\bigg)-2\int_{\mathbb{R}}\varphi(t)\cosh(\pi t)\mathrm{d}t$
$=-\sum_{\gamma}\widehat{\varphi}(\gamma)+\frac{1}{2\pi}\int_{\mathbb{R}}\widehat{\varphi}(t){\rm Re}\,\psi(\tfrac{1}{4}+i\tfrac{t}{2})\mathrm{d}t$
where $\psi(z)=\Gamma^{\prime}(z)/\Gamma(z)$ is the digamma function, $\Lambda(n)$ is the von Mangoldt function, $\rho=1/2+i\gamma$ are the zeros of the Dirichlet $L$-function
$L(\chi,s)=\sum_{n\geq 1}\chi(n)n^{-s}$
on the critical strip $0<{\rm Re}\,\rho<1$ and $\chi$ is a primitive, real and even Dirichlet character of modulus $N$. Since $|\psi(1/4+it/2)|\ll\log(2+|t|)$ and $\sum_{\gamma}\frac{1}{1+\gamma^{2}}<\infty$, we conclude, assuming the Riemann Hypothesis for $L(\chi,s)$, that if we let